# Deformation dynamics of an oil droplet into a crescent shape during the intermittent motion

### v2

@@ IV.2 Dynamics of droplet deformation induced by camphor release
The discussion of energetic aspects of droplet deformation indicates that a deformation occurs easier for a long time $t$ after the droplet is placed on the water surface. To verify this observation, we investigated deformations of paraffin droplets with oil red O that do not contain camphor. Of course, if such droplets are placed on the water surface they do not move and do not change their shapes. The surface tension gradient necessary for the motion was induced by touching the water surface close to the droplet with a camphor piece at time $T$ after the droplet had been placed on water. The experiment with a $20~{}\mu\mathrm{L}$ paraffin droplet containing oil red O (concentration $1.0~{}\mathrm{g/L}$) is illustrated in Fig. 10. As a camphor source, we used a $3~{}\mathrm{mm}$-diameter camphor disk located around $5~{}\mathrm{mm}$ away from the droplet boundary. After the camphor disk touched the surface, the droplet moved away from the source with deformation into a crescent shape, as shown in Fig. 10(a). We performed experiments for $T=30,180,300,600~{}\mathrm{s}$ and plotted the characteristic parameters describing the droplet shape dynamics in Fig. 10(b). The time $\Delta T=0$ corresponds to when the droplet COM speed $v$ has the maximum value. The maximum magnitude of deformation $A_{\mathrm{max}}$ increased with $T$ as shown in Fig. 10(b-2). The shape of the droplet was the most curved when $T=600~{}\mathrm{s}$, which can explain why $R_{\mathrm{max}}(T=600~{}\mathrm{s})$ was smaller than that when $T=300~{}\mathrm{s}$. Figure 10(b-3) shows a scatter plot of the times $\Delta T_{R}$ and $\Delta T_{A}$, at which the maximum values of $R$ and $A$ ($R_{\mathrm{max}}$ and $A_{\mathrm{max}}$) were observed. Both $\Delta T_{R}$ and $\Delta T_{A}$ were positive and had almost the same value. These behaviors were the same as those for the spontaneous deformation of a droplet with camphor shown in Fig. 4, though the relationship between $\Delta T_{R}$ and $\Delta T_{A}$ was slightly different from that between $\Delta t_{R}$ and $\Delta t_{A}$. This may be due to the much larger deformation in the experiments with a camphor-free droplet. When $T=30$ s, the droplet hardly deformed, as shown in Fig. 10(c), and the characteristic parameters could not be appropriately obtained. These results indicate that for longer $T$ of the order of $10$ minutes, the modification of the droplet interface becomes more efficient, leading to more pronounced droplet deformation.
@@
Figure 10: Deformation of a paraffin droplet containing oil red O (concentration $1.0~{}\mathrm{g/L}$) observed when the water surface was touched with a camphor disk after elapsed time $T$ after the droplet had been placed on the water surface. The time $\Delta T=0$ corresponds to when $v$ has the maximum value. (a) Snapshots of a droplet for $T=300~{}\mathrm{s}$. (b) The values of parameters characterizing droplet deformation. Panels (b-1) and (b-2) show the average value for $R_{\mathrm{max}}$ and $A_{\mathrm{max}}$ with the standard deviation as error bars for 3 experiments for each $T$, and (b-3) is a scatter plot of $\Delta T_{A}$ and $\Delta T_{R}$. (c) Superposed images of a droplet for $T=30~{}\mathrm{s}$ with a time difference of $2.0~{}\mathrm{s}$. The red line indicates the trajectories of the droplet COM.
@@ IV.3 Dynamics of flows around the droplet
In the experiments, the continuous motion changed to the intermittent motion as time passed as shown in Fig. 1. The droplet motion should be initiated by a local fluctuations in camphor surface concentration. Once such asymmetric concentration drives the droplet, the droplet keeps moving in the initial stage. In contrast, in the latter stage, the droplet cannot keep moving but shows an intermittent motion. This should be because of the less camphor release rate due to the decrease in camphor concentration in the droplet. However, the paraffin droplet with less camphor concentration did not show intermittent motion. This implies the modification of water-oil interface by the presence of oil red O molecules. The release rate may be affected by the flow around the droplet[45]. For the clarification of the more detailed mechanism of this droplet deformation phenomenon, we need more precise measurement of the dynamics of camphor and oil red O molecules. We also need to construct a mathematical model that reproduces the time evolution of motion and deformation including hydrodynamics in future study.
@@ A.1 Shape of a droplet
where $\beta=\rho g/\gamma$. The shape of the droplet can be calculated by integrating Eqs. (11)–(13) by $s$ from $s=0$, where $\phi=0$ at $s=0$. By comparing the calculated curve with the image recorded in the experiments we obtained $b$, $\beta$, and thus,the interfacial tension $\gamma$ using $\gamma=\rho g/\beta$.
@@ A.2 Analysis method
We recorded a sequence of droplet images every 60 s and calculated the interfacial tension for each image in the experiments. The droplet shape is described in the reference frame illustrated in Fig. 12(a). The vertical direction of images obtained in an experiment can differ from the actual vertical direction since the camera can be set at a slightly rotated angle. To compensate for the effect, we rotated the images with an angle $\theta$. For appropriate $\theta$, the shape of the droplet should be symmetric with respect to the vertical line. To find such an angle, the value of:
@@
was minimized. Here, $X_{r}(z,\theta)$ $(>0)$ and $X_{l}(z,\theta)$ $(<0)$ are the $x$ coordinates of the right and left peripheries at height $z$ for the rotated image. Since the position of the camera was fixed, $\theta$ should be identical for all the images of the sequence. Thus, we adopted the average of $\theta$ obtained by minimizing $E_{\theta}$ using the data of extracted 10 images. Hereafter, we use the rotated image, and the top point of the droplet estimated from the image is set to be the origin.
Next, the transformed data are compared with numerical results obtained from Eqs. (11)–(13). This calculation yields a set of coordinates $(x_{i},z_{i})$ along the edge of the droplet depending on the parameters $b$ and $\beta$. They also depend on the boundary values $x_{\mathrm{tp}}$ and $z_{\mathrm{tp}}$, which correspond to the coordinates of the top point.
@@ A.3 Derivation of the expression for line tension
$\mathcal{T}_{\mathrm{ao}}=$ $\int_{0}^{\infty}\mathrm{d}x\biggr{\{}\gamma_{\mathrm{ao}}\left(\frac{1}{\cos\theta_{\mathrm{ao}}(x)}-1\right)$
$\qquad+\frac{1}{2}\rho_{\mathrm{ao}}g(e_{\mathrm{ao}}-\zeta_{\mathrm{ao}}(x))^{2}\biggr{\}},$ (17)
$\mathcal{T}_{\mathrm{ow}}=$ $\int_{0}^{\infty}\mathrm{d}x\biggr{\{}\gamma_{\mathrm{ow}}\left(\frac{1}{\cos\theta_{\mathrm{ow}}(x)}-1\right)$
$\qquad+\frac{1}{2}\rho_{\mathrm{ow}}g(e_{\mathrm{ow}}-\zeta_{\mathrm{ow}}(x))^{2}\biggr{\}},$ (18)
@@
$\mathcal{T}_{\mathrm{ao}}=\frac{4}{3}\gamma_{\mathrm{ao}}\kappa_{\mathrm{ao}}^{-1}\left(1-\cos^{2}\frac{\theta_{\mathrm{ao}}(x=0)}{2}\right),$ (21)
$\mathcal{T}_{\mathrm{ow}}=\frac{4}{3}\gamma_{\mathrm{ow}}\kappa_{\mathrm{ow}}^{-1}\left(1-\cos^{2}\frac{\theta_{\mathrm{ow}}(x=0)}{2}\right),$ (22)

### v3

@@ IV.2 Dynamics of droplet deformation induced by camphor release
The discussion of energetic aspects of droplet deformation indicates that a deformation occurs easier for a long time $t$ after the droplet is placed on the water surface. To verify this observation, we investigated deformations of paraffin droplets with oil red O that do not contain camphor. Of course, if such droplets are placed on the water surface they do not move and do not change their shapes. The surface tension gradient necessary for the motion was induced by touching the water surface close to the droplet with a camphor piece at time $T$ after the droplet had been placed on water. The experiment with a $20~{}\mu\mathrm{L}$ paraffin droplet containing oil red O (concentration $1.0~{}\mathrm{g/L}$) is illustrated in Fig. 10. As a camphor source, we used a $3~{}\mathrm{mm}$-diameter camphor disk located around $5~{}\mathrm{mm}$ away from the droplet boundary. After the camphor disk touched the surface, the droplet moved away from the source with deformation into a crescent shape, as shown in Fig. 10(a). We performed experiments for $T=30,180,300,600~{}\mathrm{s}$ and plotted the characteristic parameters describing the droplet shape dynamics in Fig. 10(b). The time $\Delta T=0$ corresponds to when the droplet COM speed $v$ has the maximum value. The maximum magnitude of deformation $A_{\mathrm{max}}$ increased with $T$ as shown in Fig. 10(b-2). The shape of the droplet was the most curved when $T=600~{}\mathrm{s}$, which can explain why $R_{\mathrm{max}}(T=600~{}\mathrm{s})$ was smaller than that when $T=300~{}\mathrm{s}$. Figure 10(b-3) shows a scatter plot of the times $\Delta T_{R}$ and $\Delta T_{A}$, at which the maximum values of $R$ and $A$ ($R_{\mathrm{max}}$ and $A_{\mathrm{max}}$) were observed. Both $\Delta T_{R}$ and $\Delta T_{A}$ were positive and had almost the same value. This result is similar to those for the spontaneous deformation of a droplet with camphor shown in Fig. 4, though the relationship between $\Delta T_{R}$ and $\Delta T_{A}$ was slightly different from that between $\Delta t_{R}$ and $\Delta t_{A}$. However, the times $\Delta T_{R}$ and $\Delta T_{A}$ were over twice as large as $\Delta t_{R}$ and $\Delta t_{A}$. The difference can be attributed to the larger surface concentration gradient in the case of a camphor-free droplet than the one appearing for the paraffin droplet with camphor. When $T=30$ s, the droplet hardly deformed, as shown in Fig. 10(c), and the characteristic parameters could not be appropriately obtained. These results indicate that for longer $T$ of the order of $10$ minutes, the modification of the droplet interface becomes more efficient, leading to more pronounced droplet deformation.
@@
Figure 10: Deformation of a paraffin droplet containing oil red O (concentration $1.0~{}\mathrm{g/L}$) observed when the water surface was touched with a camphor disk at time $T$ after the droplet had been placed on the water surface. The time $\Delta T=0$ corresponds to when $v$ has the maximum value. (a) Snapshots of a droplet for $T=300~{}\mathrm{s}$. (b) The values of parameters characterizing droplet deformation. Panels (b-1) and (b-2) show the average value for $R_{\mathrm{max}}$ and $A_{\mathrm{max}}$ with the standard deviation as error bars for 3 experiments for each $T$, and (b-3) is a scatter plot of $\Delta T_{A}$ and $\Delta T_{R}$. (c) Superposed images of a droplet for $T=30~{}\mathrm{s}$ with a time difference of $2.0~{}\mathrm{s}$. The red line indicates the trajectories of the droplet COM.
@@ IV.3 Dynamics of flows around the droplet
In the experiments, the continuous motion changed to the intermittent motion circa half an hour after a droplet was placed on the water surface as shown in Fig. 1. In the continuous motion, asymmetric concentration driving the droplet is supported by coupling between motion and high rate of camphor release. The transition to intermittent motion can be explained by combination of two factors. One is a decrease in camphor release rate due to declining camphor concentration in the droplet. However, the decrease alone does not fully explain the transition because paraffin droplets with small camphor concentration did not show intermittent motion. Our analysis shows that the modification of the water-oil interface by the presence of oil red O molecules is the factor allowing for moderation of camphor release as well as for changes in droplet shape. The release rate may be additionally affected by the flow around the droplet[46]. For the clarification of the more detailed mechanism of this droplet deformation phenomenon, we need more precise measurement of the dynamics of camphor and oil red O molecules. We also need to construct a mathematical model that reproduces the time evolution of motion and deformation including hydrodynamics in future study.
@@ A.1 Shape of a droplet
where $\beta=\rho g/\gamma$. The shape of the droplet can be calculated by integrating Eqs. (9)–(11) by $s$ from $s=0$ with $\phi(s=0)=0$. By comparing the calculated curve with the image recorded in the experiments we obtained $b$, $\beta$, and thus,the interfacial tension $\gamma$ using $\gamma=\rho g/\beta$.
@@ A.2 Analysis method
We recorded a sequence of droplet images every 60 s and calculated the interfacial tension for each image in the experiments. The droplet shape is described in the reference frame illustrated in Fig. 12(a). The vertical direction of images obtained in an experiment can differ from the actual vertical direction since the camera can be set at a slightly rotated angle. To compensate for the effect, we rotated the images with an angle $\theta$ for the most symmetric distribution of the image with respect to the vertical line. To find $\theta$, the value of:
@@
was minimized. Here, $X_{r}(z,\theta)$ $(>0)$ and $X_{l}(z,\theta)$ $(<0)$ are the $x$ coordinates of the right and left peripheries at height $z$ for the rotated image. Since the position of the camera was fixed, $\theta$ should be identical for all the images of the sequence. Thus, we adopted the average of $\theta$ obtained by minimizing $E_{\theta}$ using the data of extracted 10 images. In the following, we use the rotated image, and the top point of the droplet estimated from the image is set to be the coordinate origin.
Next, the transformed data are compared with numerical results obtained from Eqs. (9)–(11). This calculation yields a set of coordinates $(x_{i},z_{i})$ along the edge of the droplet depending on the parameters $b$ and $\beta$ as well as on the coordinates of the top point $x_{\mathrm{tp}}$ and $z_{\mathrm{tp}}$.
@@ A.3 Derivation of the expression for line tension
$\mathcal{T}_{\mathrm{ao}}=$ $\int_{0}^{\infty}\mathrm{d}x\left[\gamma_{\mathrm{ao}}\left(\frac{1}{\cos\theta_{\mathrm{ao}}(x)}-1\right)\right.$
$\qquad\left.+\frac{1}{2}\rho_{\mathrm{ao}}g(e_{\mathrm{ao}}-\zeta_{\mathrm{ao}}(x))^{2}\right],$ (15)
$\mathcal{T}_{\mathrm{ow}}=$ $\int_{0}^{\infty}\mathrm{d}x\left[\gamma_{\mathrm{ow}}\left(\frac{1}{\cos\theta_{\mathrm{ow}}(x)}-1\right)\right.$
$\qquad+\left.\frac{1}{2}\rho_{\mathrm{ow}}g(e_{\mathrm{ow}}-\zeta_{\mathrm{ow}}(x))^{2}\right],$ (16)
@@
$\mathcal{T}_{\mathrm{ao}}=\frac{4}{3}\gamma_{\mathrm{ao}}\kappa_{\mathrm{ao}}^{-1}\left[1-\cos^{2}\left(\frac{\theta_{\mathrm{ao}}(x=0)}{2}\right)\right],$ (19)
$\mathcal{T}_{\mathrm{ow}}=\frac{4}{3}\gamma_{\mathrm{ow}}\kappa_{\mathrm{ow}}^{-1}\left[1-\cos^{2}\left(\frac{\theta_{\mathrm{ow}}(x=0)}{2}\right)\right],$ (20)
@@
Assuming that the water surface outside of the droplet is completely flat, we obtain
$\cos\left(\theta_{\mathrm{ao}}(x=0)\right)$ $=\frac{\gamma_{\mathrm{aw}}^{2}-\gamma_{\mathrm{ow}}^{2}+\gamma_{\mathrm{ao}}^{2}}{2\gamma_{\mathrm{aw}}\gamma_{\mathrm{ao}}},$ (21)
$\cos\left(\theta_{\mathrm{ow}}(x=0)\right)$ $=\frac{\gamma_{\mathrm{aw}}^{2}-\gamma_{\mathrm{ao}}^{2}+\gamma_{\mathrm{ow}}^{2}}{2\gamma_{\mathrm{aw}}\gamma_{\mathrm{ow}}},$ (22)
which relates to the excess energy of the triple line with the values of interfacial tensions.